OpenStudy (anonymous):

y^3+6y^2+9y,y^2-6y find the LCM

OpenStudy (anonymous):

Upgrade level? A lil more challenging :) y^3 +6y^2 + 9y= y(y^2 +6y+9)=y(y+3)(y+3) y^2 -6y= y(y-6) LCD= y(y+3)(y+3)(y-6)